What is the future value of $4,000 in a bank account for 7 years at 5 percent compounded bimonthly?
Answer by Alan3354(69443)

What is the future value of $4,000 in a bank account for 7 years at 5 percent compounded bimonthly?
-----------------
FV = 4000*(1 + r)^t
r = 0.05/24
t = 24*7 = 168